OpenStudy (anonymous):

One of your friends sends you an email asking you to explain how all of the following expressions have the same answer. the cube root of x cubed :x to the one–third power • :x to the one–third power • x to the one–third power 1 over x to the –1 power the eleventh root of the quantity of x to the fifth times x to the fourth times x squared

OpenStudy (akashdeepdeb):

\[\sqrt[3]{x^3} = x\] \[x^{1/3} * x^{1/3} * x^{1/3} = x^{\frac{ 1 }{ 3 } + \frac{ 1 }{ 3 } + \frac{ 1 }{ 3 } } = x^{3/3} = x^1 = x\] \[(\frac{ 1 }{ x })^{-1} = x^1 = x\] \[\sqrt[11]{x^5 * x^4 * x^2} = \sqrt[11]{x^{5+4+2}} = \sqrt[11]{x^{11}} = x^{11/11} = x^1 = x \] Understood why they are the same? :)

OpenStudy (anonymous):

Yes, I can see now how they are related to each other when they are put like that.

OpenStudy (anonymous):

Thank you for your help :)

OpenStudy (akashdeepdeb):

:)